Using the appropriate interest table, compute the present values of the following periodic amounts due at the end of the designated periods.

Collapse question part

(a)

Correct answer. Your answer is correct.
Click here to view factor tables

$52,710 receivable at the end of each period for 9 periods compounded at 12%. (Round factor values to 5 decimal places, e.g. 1.25124 and final answer to 0 decimal places, e.g. 458,581.)
Present value $enter the present value in dollars rounded to 0 decimal places

Entry field with correct answer280852

SHOW SOLUTION

SHOW ANSWER

LINK TO TEXT

By accessing this Question Assistance, you will learn while you earn points based on the Point Potential Policy set by your instructor.
Attempts: 1 of 3 used

Collapse question part

(b)

Click here to view factor tables

$52,710 payments to be made at the end of each period for 17 periods at 9%. (Round factor values to 5 decimal places, e.g. 1.25124 and final answer to 0 decimal places, e.g. 458,581.)
Present value $enter the present value in dollars rounded to 0 decimal places
Answer #1

a. Present value = 52710*5.3282 = $280852

b. Present value = 52710*8.5436 = $450333